### v2

# Periodic Vibration Gaussian: Dynamic Urban Scene Reconstruction and Real-time Rendering

Abstract
Modeling dynamic, large-scale urban scenes is challenging due to their highly intricate geometric structures and unconstrained dynamics in both space and time. Prior methods often employ high-level architectural priors, separating static and dynamic elements, resulting in suboptimal capture of their synergistic interactions. To address this challenge, we present a unified representation model, called Periodic Vibration Gaussian (PVG). PVG builds upon the efficient 3D Gaussian splatting technique, originally designed for static scene representation, by introducing periodic vibration-based temporal dynamics. This innovation enables PVG to elegantly and uniformly represent the characteristics of various objects and elements in dynamic urban scenes. To enhance temporally coherent and large scene representation learning with sparse training data, we introduce a novel temporal smoothing mechanism and a position-aware adaptive control strategy respectively. Extensive experiments on Waymo Open Dataset [31] and KITTI benchmarks [13] demonstrate that PVG surpasses state-of-the-art alternatives in both reconstruction and novel view synthesis for both dynamic and static scenes. Notably, PVG achieves this without relying on manually labeled object bounding boxes or expensive optical flow estimation. Moreover, PVG exhibits 900-fold acceleration in rendering over the best alternative.
(a) Dynamic scene
(b) Remove the dynamic
Figure 1: Our proposed Periodic Vibration Gaussian is crafted to effectively and uniformly capture both the static and dynamic elements of a large, dynamic urban scene. (a) It not only reconstructs a dynamic urban scene but also enables real-time rendering, while efficiently isolating dynamic components from the intricacies of the highly unconstrained and complex scene. (b) This capability facilitates flexible manipulation, such as the removal of dynamic scene elements.
1 Introduction
The geometric reconstruction of extensive urban spaces, such as streets and cities, has played a pivotal role in applications like digital maps, auto-navigation, and autonomous driving [31, 13, 6]. Our world is inherently dynamic and complex in both spatial and temporal dimensions. Despite advancements in scene representation techniques like Neural Radiance Fields (NeRFs) [28, 43, 14], which primarily focus on static scenes, they overlook more challenging dynamic elements.
Recent approaches to model dynamic scenes include NSG [23], which decomposes dynamic scenes into scene graphs and learns a structured representation. PNF [17] further decomposes scenes into objects and backgrounds, incorporating a panoptic segmentation auxiliary task. However, scalability issues arise in real-world scenarios, where obtaining accurate object-level supervisions (e.g., 3D object boxes, segmentation masks) is challenging, and explicitly representing each object linearly increases model complexity with the number of objects.
SUDS [35] later proposes using optical flow to relax the stringent requirement of object labeling in a three-branch architecture for separately modeling static and dynamic elements and the environmental factors of a scene. EmerNeRF [45] uses a self-supervised method to reduce dependence on optical flow. Despite adopting implicit NeRF representation, these methods suffer from low efficiency in both training and rendering, posing a significant bottleneck for large-scale scene rendering and reconstruction. Additionally, manually separating constituent parts introduces design complexity and limits the ability to capture intrinsic correlations and interactions.
To overcome the identified limitations, this paper introduces a novel dynamic scene representation method termed Periodic Vibration Gaussian (PVG). This approach provides a unified representation of both static and dynamic elements within a scene through a single formulation. Building upon the efficient 3D Gaussian Splatting technique [15], originally devised for static scene representation, we incorporate periodic vibration-based temporal dynamics. This modification allows for a cohesive representation of static and dynamic scene elements with explicit motion properties such as velocity and staticness. We also propose a position-aware point adaptive control strategy to fit the distant view better. To improve the temporal continuity in representation learning with typically limited training data, we introduce a novel temporal smoothing mechanism.
Our contributions are summarized as follows: (i) Introduction of the very first unified representation model, Periodic Vibration Gaussian (PVG), for large-scale dynamic urban scene reconstruction. In contrast to previous NeRF-based solutions, PVG employs the 3D Gaussian Splatting paradigm, uniquely extending it to elegantly represent dynamic scenes. This is accomplished by seamlessly integrating periodic vibration-based temporal dynamics into the conventional 3D Gaussian Splatting formulation. (ii) Development of a novel temporal smoothing mechanism to enhance the temporal continuity of representation and a position-aware point adaptive control strategy for unbounded urban scenes. (iii) Extensive experiments on two large benchmarks (KITTI and Waymo) demonstrate that PVG outperforms all previous state-of-the-art alternatives in novel view synthesis. Moreover, it provides significant efficiency benefits in both training and rendering processes, achieving a remarkable 900-fold acceleration in rendering compared to the leading competitor, EmerNeRF [45].
2 Related work
Neural rendering In the domain of novel view synthesis, Neural Radiance Fields (NeRF) [21] have emerged as a noteworthy approach. NeRF employs a coordinate-based multi-layer perception representation of 3D scenes, leveraging volumetric rendering and the spatial smoothness of multi-layer perception to generate high-quality novel views. However, its implicit nature comes with significant drawbacks, including slow training and rendering speeds, as well as high memory usage.
To tackle these challenges, several studies have proposed solutions to enhance training speed. Techniques such as voxel grids [30], hash encoding [22], and tensor factorization [8, 9] have been explored. Others have focused on improving rendering speed by transforming implicit volumes into explicit textured meshes, as demonstrated in works like [10, 27, 49]. Additionally, endeavors such as [3, 4, 36, 5] aim to enhance rendering quality by addressing issues like antialiasing and reflection modeling. More recently, 3D Gaussian Splatting (3DGS) introduced an innovative point-based 3D scene representation, seamlessly integrating the high-quality volume rendering principles of NeRF [21] with the swift rendering speed characteristic of rasterization.
Dynamic scene models Reconstructing dynamic scenes poses distinctive challenges, particularly in effectively handling temporal correlations across various time steps. Expanding on the accomplishments of NeRF [21], several extensions have been proposed to tailor NeRF to dynamic scenarios. In one research direction, certain studies [18, 11, 7, 38, 2] introduce time as an additional input to the radiance field, treating the scene as a 6D plenoptic function. However, this approach couples positional variations induced by temporal dynamics with the radiance field, lacking geometric priors about how time influences the scene. An alternative approach [26, 24, 25, 33, 1, 20, 40] focuses on modeling the movement or deformation of specific static structures, assuming that the dynamics arise from these static elements within the scene. Point-based methods [1, 20, 40] have shown promise in addressing the challenges of reconstructing dynamic scenes due to their flexibility. Building upon the progress in 3D Gaussian splatting, recent works [48, 40] propose the use of a set of deformable 3D Gaussians for optimization across different timestamps. While this kind of methods need to learn the deform function in the dense space which is difficult to migrate to large scenarios. [47] extends the 3D gaussian expression to a 4D form leading to confusion between dynamic and static distinctions.
Urban scene reconstruction NeRF-based techniques have shown their efficacy in autonomous driving scenarios [13, 31]. One research avenue has focused on enhancing the modeling of static street scenes by utilizing scalable representations [32, 34, 19, 28], achieving high-fidelity surface reconstruction [28, 39, 14], and incorporating multi-object composition [43]. However, these methods face difficulties in handling dynamic elements commonly encountered in autonomous driving contexts. Another research direction seeks to address these challenges. Notably, these techniques require additional input, such as leveraging panoptic segmentation to refine the dynamics of reconstruction [17]. Moreover, in [23], scene graphs are employed to decompose dynamic multi-object scenes, while in [46], neural shape priors are learned for completing dynamic object reconstructions. In [41] foreground instances and background environments are decomposed. [44, 50] decompose the scene with different set of gaussian points by bounding boxes. However, they all need manually annotated or predicted bounding boxes and have difficulty to reconstruct the non-rigid objects. In [35], a scalable hash table is proposed for large-scale dynamic scenes, relying on an off-the-shelf 2D optical flow estimator to track dynamic actors. [45] reduces dependence on optical flow by self-supervision, however, it still suffers low image quality and rendering speed.
In this paper, we present an elegant extension of 3D Gaussian Splatting [15] with the additional time dimension to handle the complexities of dynamic scenes. Our model provides a uniform, efficient representation, excelling in reconstructing dynamic, large-scale urban scenes without the dependence on manual annotations or pre-trained models.
3 Method
Utilizing the sequentially acquired and calibrated multi-sensor data, encompassing a set of images $\mathcal{I}$, each captured by cameras equipped with corresponding intrinsic matrices $\mathbf{I}$ and extrinsic matrices $\mathbf{E}$, alongside their respective capture timestamps $t$, collectively represented as $\{\mathcal{I}_{i},t_{i},\mathbf{E}_{i},\mathbf{I}_{i}|i=1,2,\dots N_{c}\}$, and the spatial coordinates of LiDAR point clouds annotated with timestamps $\{(x_{i},y_{i},z_{i},t_{i})|i=1,2,\dots N_{l}\}$, where $N_{c}$ and $N_{l}$ are the number of image frames and Lidar points, our research aims to achieve precise 3D reconstruction and synthesize novel viewpoints at any desired timestamp $t$ and camera pose $[\mathbf{E}_{o},\mathbf{I}_{o}]$. To this end, our framework is meticulously engineered to approximate a rendering function ^$\hat{\mathcal{I}}=\mathcal{F}_{\theta}(\mathbf{E}_{o},\mathbf{I}_{o},t)$.
3.1 Preliminary
3D Gaussian Splatting (3DGS) [15] utilizes a collection of 3D Gaussians to represent a scene. Through a tile-based rasterization process, 3DGS facilitates real-time alpha blending of numerous Gaussians. The scene is modeled by a set of points $\{P_{i}\}$, where each point $P$ is linked to a mean $\bm{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, an opacity $o$, and a color $\mathbf{c}$. These attributes collectively define the point’s influence within the 3D space as:
$G(\bm{x})~{}=e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\Sigma^{-1}(\bm{x}-\bm{\mu})}.$ (1)
To create an image from a particular viewpoint, 3DGS maps each Gaussian point onto the image plane, yielding a collection of 2D Gaussians. Calculating the means of these projected Gaussians is straightforward. However, determining the formulation for the projected variance involves
$\Sigma^{\prime}=JW~{}\Sigma~{}W^{T}J^{T},$ (2)
where $W$ and $J$ are the view transform matrix and Jacobian of the nonlinear projective transform matrix, respectively. Sorting the Gaussians according to their depth in camera space, we can query the attributes of each 2D Gaussian and facilitate the subsequent volume rendering process to determine the color of each pixel:
$C=\sum^{N}_{i=1}T_{i}\alpha_{i}\bm{c}_{i}\hskip 5.0pt\\ \text{ with }\hskip 5.0ptT_{i}=\prod^{i-1}_{j=1}(1-\alpha_{j}),$ (3)
where $\alpha$ is derived through the product of the opacity ${o}$ and the contribution from the 2D covariance calculated using $\Sigma^{\prime}$ and the corresponding pixel coordinates in the image space. The covariance matrix holds a meaningful interpretation when it is positive semi-definite. In the context of 3DGS, it is decomposed into a scaling matrix, denoted as a diagonal matrix represented by $\bm{s}\in\mathbb{R}^{3}$, and a rotation matrix represented by a unit quaternion $\bm{q}$.
The differentiable rendering function for a new view of a scene containing $N$ points is expressed as
^$\hat{\mathcal{I}}=\mathrm{Render}(\{\mathcal{C}_{i}|i=1,2,\dots,N\};\mathbf{E},\mathbf{I}),$ (4)
where $\hat{\mathcal{I}}$ is the rendered image, and $\mathbf{E}$ and $\mathbf{I}$ denote the camera extrinsic and intrinsic matrices, respectively. The training of the model entails optimizing the parameter set for each point, represented as $\mathcal{C}=\{\bm{\mu},\bm{q},\bm{s},o,\bm{c}\}$.
Flexible rendering This rendering method can tackle different targets, like depth and opacity, by replacing the color $\bm{c}$ in Eq. (3). For example, the normalized depth map can be computed as: $\sum^{N}_{i=1}T_{i}\alpha_{i}z_{i}/\sum^{N}_{i=1}T_{i}\alpha_{i}$, where $z_{i}$ represents the distance of the center of a Gaussian point from the image plane.
Limitation The 3D Gaussian Splatting model represents static points in a scene, lacking the ability to capture dynamic changes over time, essential for modeling dynamic urban scenes. To address this limitation, we propose the Periodic Vibration Gaussian model.
3.2 Periodic Vibration Gaussian (PVG)
Figure 2: (a) Dynamic and static can be expressed uniformly through different Periodic Vibration Gaussian points with different lifespan. Our model can be trained to differentiate between dynamic and static on its own. (b) Our training pipeline.
Our PVG model exhibits several distinctive features:
Dynamics introduction: We introduce the concept of life peak, denoted as $\tau$, which represents the point’s moment of maximum prominence over time. The motivation behind this concept is to assign a distinct lifespan to each Gaussian point, defining when it actively contributes and to what degree. This fundamentally infuses a dynamic nature into the model, enabling variations in the collection of Gaussian points that influence the rendering of the scene over time.
Periodic vibration: We modify the traditional 3D Gaussian’s mean $\bm{\mu}$ and opacity $o$ to be time-dependent functions centered around the life peak $\tau$, denoted as $\widetilde{\bm{\mu}}(t)$ and $\widetilde{o}(t)$. Both functions peak at $\tau$. This adaptation empowers the model to effectively capture dynamic motions, enabling each point to adjust based on temporal changes.
Formally, our model, denoted as $\mathcal{H}$, is expressed as:
$\mathcal{H}(t)$ =$=\{\widetilde{\bm{\mu}}(t),\bm{q},\bm{s},\widetilde{o}(t),\bm{c}\},$ (5)
$\quad\widetilde{\bm{\mu}}(t)$ $=\bm{\mu}+\frac{l}{2\pi}\cdot\sin(\frac{2\pi(t-\tau)}{l})\cdot\bm{v},$ (6)
$\widetilde{o}(t)$ $=o\cdot e^{-\frac{1}{2}(t-\tau)^{2}\beta^{-2}},$ (7)
where $\widetilde{\bm{\mu}}(t)$ represents the vibrating motion centered at $\bm{\mu}$ occurring at the life peak $\tau$, and $\widetilde{o}(t)$ denotes the vibrating opacity, which decays away from the peak $\tau$ with a decay rate inversely proportional to $\beta$. Notably, the parameter $\beta$ governs the lifespan around $\tau$, with larger values indicating bigger lifespans. The hyper-parameter $l$ represents the cycle length, serving as the scene prior. The learnable parameter $\bm{v}=\frac{\mathrm{d}\widetilde{\bm{\mu}}(t)}{\mathrm{d}t}|_{t=\tau}\in\mathbb{R}^{3}$ signifies the vibrating direction and denotes the instant velocity at time $\tau$. Therefore, the per-point learnable parameters of our model $\mathcal{H}$ include $\{\bm{\mu},\bm{q},\bm{s},o,\bm{c},\tau,\beta,\bm{v}\}$.
In particular, we express the mean vector (position) $\widetilde{\bm{\mu}}(t)$ through periodic vibrations, providing a cohesive framework for both static and dynamic patterns. To enhance clarity, we introduce the staticness coefficient $\rho=\frac{\beta}{l}$, which quantifies the degree of staticness exhibited by a PVG point and is also associated with the point’s lifespan. Periodic vibration facilitates convergence around $\bm{\mu}$ when $\rho$ is large. This is due to the bounded nature of $\widetilde{\bm{\mu}}(t)$ by $\bm{v}$ and the fact that $\mathbb{E}[\widetilde{\bm{\mu}}(t)]=\bm{\mu}$ holds for any time interval with a length that is a multiple of $l$, independent of $\bm{v}$.
We note that the 3DGS represents a particular case of PVG in which $\bm{v}=\mathbf{0}$ and $\rho=+\infty$. PVGs with large $\rho$ effectively capture the static aspects in a scene, provided that $\|\bm{v}\|$ remains within a reasonable range.
The ability of our PVG to represent the dynamic aspects of a scene is particularly evident in points with small $\rho$. Points approaching $\rho\to 0$ manifest by appearing and disappearing almost instantaneously, executing linear movements around the time $\tau$. As time progresses, these points undergo oscillations, with some vanishing and others emerging. At a specific timestamp $t$, dynamic objects are more likely to be predominantly represented by points with $\tau$ close to $t$. In essence, different points take charge of representing dynamic objects at distinct timestamps.
Conversely, the static components of a scene can be effectively represented by points exhibiting large $\rho$. Introducing a threshold on $\rho$ enables us to discern whether a point represents dynamic elements (see Fig. 1(b)).
It is crucial to emphasize that at any given time $t$, our model takes the form of a specific 3D Gaussian model, represented by $\mathcal{H}(t)$. We train a collection of PVG points, denoted as $\{\mathcal{H}_{i}\}$, to effectively portray a dynamic scene. The rendering process is then executed as:
^$\hat{\mathcal{I}}=\mathrm{Render}(\{\mathcal{H}_{i}(t)|i=1,\dots,N_{H}\};\mathbf{E},\mathbf{I}),$ (8)
where $N_{H}$ represents the number of PVG points in a scene. Our training pipeline can be seen in Fig. 2(b).
3.3 Position-aware point adaptive control
The conventional adaptive control method, as introduced in [15], treating each Gaussian point uniformly, proves inadequate for urban scenes. This is mainly attributed to the substantial distance of the mean vector (position) $\bm{\mu}$ for most points from the center of the unbounded scene. To faithfully represent the scene with fewer points without sacrificing accuracy, we advocate utilizing larger points for distant locations and smaller points for nearby areas.
Assuming camera poses are centered, the inclusion of the scale factor $\gamma(\bm{\mu})$ as defined below is essential for effective control over each PVG point:
$\begin{split}\gamma(\bm{\mu})=\left\{\begin{array}[]{ll}1&\mathrm{if}\quad\|\bm{\mu}\|_{2}<2r\\ \|\bm{\mu}\|_{2}/r-1&\mathrm{if}\quad\|\bm{\mu}\|_{2}\geq 2r,\end{array}\right.\end{split}$ (9)
where $r$ denotes the scene radius (i.e., the scene scope). Specifically, we employ a densification strategy for a PVG $\mathcal{H}(t)$ when its backward gradient on view space surpasses a specified threshold. We opt to clone the PVG if $\max(\bm{s})\leq g\cdot\gamma(\bm{\mu})$, with $g$ serving as the threshold for scale. Conversely, if this condition is not met, we initiate a split operation. Additionally, we undertake pruning of points with $\max(\bm{s})>b\cdot\gamma(\bm{\mu})$, employing $b$ as the scale threshold to discern whether a given PVG is excessively large.
3.4 Model training
Figure 3: Our temporal smoothing mechanism. We query the status of PVG point set at $t-\Delta t$ and translate each point with its 3D flow translation $\bm{\bar{v}}\cdot\Delta t$, we further render the translated set of points to do the supervision at timestamp $t$.
Temporal smoothing by intrinsic motion Reconstructing dynamic scenes in autonomous driving poses a significant challenge, primarily attributed to the sparse data in terms of both views and timestamps, as well as the unconstrained variations across frames. Specifically, in PVG, individual points encompass only a narrow time window, resulting in constrained training data and an increased susceptibility to overfitting. To address this, we capitalize on the inherent dynamic properties of PVG, which establish connections between the states of consecutive observations.
Instead of flow estimation, we introduce the average velocity metric as:
$\bm{\bar{v}}=\frac{\mathrm{d}\widetilde{\bm{\mu}}(t)}{\mathrm{d}t}\bigg{|}_{t=\tau}\cdot\exp(-\frac{\rho}{2})=\bm{v}\cdot\exp(-\frac{\rho}{2}).$ (10)
The intuition comes from the average velocity weighted by opacity decay which is hard to compute the explicit solution, i.e.,
$\bm{\bar{v}}\propto\frac{1}{\sqrt{2\pi}\beta}\int^{+\infty}_{-\infty}\frac{\mathrm{d}\widetilde{\bm{\mu}}(t)}{\mathrm{d}t}\cdot e^{-\frac{1}{2}(t-\tau)^{2}\beta^{-2}}\mathrm{d}t.$ (11)
This metric is bounded as it satisfies that $\lim_{\rho\to\infty}\bm{\bar{v}}=\mathbf{0}$, and $\lim_{\rho\to 0}\bm{\bar{v}}=\bm{v}$.
In practical scenarios, dynamic objects often maintain a constant speed within a short time interval. This observation leads to the emergence of a linear relationship between consecutive states of PVG.
Formally, consider two adjacent timestamps, $t_{1}$ and $t_{2}$ ($t_{1}<t_{2}$) with their respective states represented as $\{\mathcal{H}_{i}(t_{1})\}$ and $\{\mathcal{H}_{i}(t_{2})\}$. These states are linearly connected by a flow translation for each point, denoted as $\Delta\bm{\mu}=\bm{\bar{v}}\cdot(t_{2}-t_{1})=\bm{\bar{v}}\Delta t$. Specifically, we estimate the underlying state of ${\mathcal{H}}(t_{2})$ as:
^$\widehat{\mathcal{H}}(t_{2})=\{\widetilde{\bm{\mu}}(t_{1})+\bm{\bar{v}}\cdot\Delta t,\bm{q},\bm{s},\widetilde{o}(t_{1}),\bm{c}\}.$ (12)
We note that this estimation process is applied to each individual PVG point. A visual representation of this estimation is illustrated in Fig. 3.
We utilize estimated states to improve model training. Specifically, we assign a probability of $\eta$ to set $\Delta t$ as 0 (indicating no estimation), and for the remaining probability, we randomly sample $\Delta t$ from a uniform distribution $\mathrm{U}(-\delta,+\delta)$. In the latter case, we replace $\mathcal{H}$ (Eq. (8)) with $\widehat{\mathcal{H}}$ during training. This strategy helps each point to learn its correct motion trend from the adjacent training frames which acts like a self-supervision mechanism, fostering a more consistent representation without imposing a significant increase in computational demands as well as allowing eliminate the dependence on optical flow estimation. By adopting this approach, we improve temporal coherence and consistency, thereby alleviating the challenges posed by sparse data and the risk of overfitting.
Sky refinement To handle the high-frequency details in the sky, we utilize a high-resolution learnable environment cube map $f_{sky}(d)=c_{sky}$ as the background. The final color is articulated as $C_{f}=C+(1-O)f_{sky}(d)$, where $O=\sum^{N_{H}}_{i=1}T_{i}\alpha_{i}$ represents the rendered opacity. During the training phase, we incorporate random perturbations to the ray direction $d$ within its unit pixel length to enhance anti-aliasing.
Objective Our objective loss function is formulated as:
$\mathcal{L}=(1-\lambda_{r})\mathcal{L}_{1}+\lambda_{r}\mathcal{L}_{\mathrm{ssim}}+\lambda_{d}\mathcal{L}_{d}+\lambda_{o}\mathcal{L}_{o}+\lambda_{\bm{\bar{v}}}\mathcal{L}_{\bm{\bar{v}}},$ (13)
where $\mathcal{L}_{1}$ and $\mathcal{L}_{\mathrm{ssim}}$ are L1 and SSIM loss [15] for supervision of RGB rendering.
The term $\mathcal{L}_{d}=\frac{1}{hw}\sum\|\mathcal{D}^{s}-\mathcal{D}\|_{1}$ is a depth loss for geometry awareness, where $\mathcal{D}^{s}$ is a sparse inverse depth map generated by projecting the LiDAR points to the camera plane, $\mathcal{D}$ denotes the inverse of the rendered depth map, and $h$ and $w$ denote the rendering spatial size.
The term $\mathcal{L}_{o}=-\frac{1}{hw}\sum O\cdot\log O-\frac{1}{hw}\sum M_{sky}\cdot\log(1-O)$ is the opacity loss where $M_{sky}$ is the sky mask estimated by a pretrained segmentation model [42]. This loss aims to drive the opacity values towards either 0 (representing a transparent sky) or 1. Specifically, it regularizes opacity to 0 for predicted sky pixels.
The last term $\mathcal{L}_{\bm{\bar{v}}}=\frac{1}{hw}\sum\|\mathcal{\bar{V}}\|_{1}$ is the sparse velocity loss where $\mathcal{\bar{V}}$ is the rendered average velocity $\bm{\bar{v}}$ map. This loss not only leads to a sparse $\|\bm{v}\|$ but also encourages larger $\beta$ (corresponding to static scene components). The rationale behind this is that most elements of a scene are static.
4 Experiments
(a) 3DGS [15]
(b) SUDS [35]
(c) EmerNeRF [45]
(d) PVG (Ours)
(e) GT
Figure 4: Novel view synthesis of dynamic scenes on Waymo.
(a) 3DGS [15]
(b) SUDS [35]
(c) EmerNeRF [45]
(d) PVG (Ours)
(e) GT
Figure 5: Novel view synthesis of dynamic scenes on KITTI.
Competitors For dynamic scenes, we evaluate our method alongside S-NeRF [43], StreetSurf [14], Mars [41], 3DGS [15], NSG [23], SUDS [35] and EmerNeRF [45]. In line with previous methods, we conduct evaluations on both image reconstruction and novel view synthesis tasks, selecting every fourth timestamp from each camera as the test set for novel view synthesis. Although our primary focus is on dynamic scenes, to be fair, we also provide a quantitative comparison with methods [43, 14, 15] tailored for static scenes. This is to demonstrate our model’s superiority at uniformly managing both static and dynamic environments.
Implementation details For points initialization, we sample $6\times 10^{5}$ LiDAR points, $2\times 10^{5}$ near points whose distance to the origin is uniformly sampled from $(0,r)$, $2\times 10^{5}$ far points whose inverse distance is uniformly sampled from $(0,1/r)$, where $r$ is the foreground radius different across scenes, around 30 meters. $\beta$ is set to $0.3$, $\bm{v}$ is set to $\mathbf{0}$. We employ the Adam optimizer [16] and maintain a similar learning rate for most parameters as the original 3DGS implementation while we adjust the learning rate of the velocity $\bm{v}$, opacity decaying $\beta$ and opacity $o$ to $1\times 10^{-3}$, $0.02$ and $0.005$ respectively. Regarding the densification schedule, we set the image space densification threshold to $1.7\times 10^{-4}$ and reset the opacity of Gaussians to $0.01$ every $3,000$ iterations to remove superfluous points. For regularization, we use coefficients $\lambda_{r}=0.2$, $\lambda_{d}=0.1$, $\lambda_{o}=0.05$ and $\lambda_{\bm{\bar{v}}}=0.01$. For Temporal smoothing training mechanism, we uniformly sample time intervals $\Delta t$ from a 1.5 frame span in the camera sequence with a probability $\eta=0.5$. We set the cube map resolution to $1024$ to capture the high-frequency details in the sky. The training process commences from a downsampled scale of $16$, which is then gradually increased every $5,000$ iterations. We conduct all experiments on a single NVIDIA RTX A6000 GPU for a total of $30,000$ iterations which takes about an hour to yield the final results, and the rendering speed can achieve 50 FPS. We rescale the time interval between two consecutive frames to 0.02 and fix $l=0.2$. As shown in Fig. 1, we remove the PVG points whose $\rho<1$ to preserve the static part of the scene.
Metrics Consistent with SUDS [35], we use the PSNR, SSIM, and LPIPS metrics to measure image reconstruction and novel view synthesis.
4.1 Comparison with state of the art
Table 1: Quantitative results on dynamic scenes on Waymo and KITTI.
Waymo Open Dataset KITTI
Image reconstruction Novel view synthesis Image reconstruction Novel view synthesis
[…]
S-NeRF [43] 0.0014 19.67 0.528 0.387 19.22 0.515 0.400 0.0075 19.23 0.664 0.193 18.71 0.606 0.352
StreetSurf [14] 0.097 26.70 0.846 0.3717 23.78 0.822 0.401 0.37 24.14 0.819 0.257 22.48 0.763 0.304
3DGS [15] 63 27.99 0.866 0.293 25.08 0.822 0.319 125 21.02 0.811 0.202 19.54 0.776 0.224
NSG [23] 0.032 24.08 0.656 0.441 21.01 0.571 0.487 0.19 26.66 0.806 0.186 21.53 0.673 0.254
Mars [41] 0.030 21.81 0.681 0.430 20.69 0.636 0.453 0.31 27.96 0.900 0.185 24.23 0.845 0.160
SUDS [35] 0.008 28.83 0.805 0.289 21.83 0.656 0.405 0.04 28.31 0.876 0.185 22.77 0.797 0.171
EmerNeRF [45] 0.053 28.11 0.786 0.373 25.92 0.763 0.384 0.28 26.95 0.828 0.218 25.24 0.801 0.237
PVG (Ours) 50 32.46 0.910 0.229 28.11 0.849 0.279 59 32.83 0.937 0.070 27.43 0.896 0.114
PVG (5cam) 41 32.05 0.895 0.252 27.50 0.828 0.312 - - - - - - -
Results on Waymo The Waymo Open Dataset encompasses over 1,000 driving segments, each with a duration of 20 seconds, recorded using five high-resolution LiDARs and five cameras facing the front and sides.
In our experiments, we utilize the three frontal cameras to assess performance on four challenging dynamical scenes (each contains around 50 frames), chosen due to their substantial movement. Table 1 displays the average error metrics across these selected dynamic scenes for both image reconstruction and novel view synthesis tasks. Our model markedly outperforms baselines [43, 14, 15, 23, 35, 41] across all metrics in both tasks. Specifically, for the image reconstruction task, we note a $12.6\%$ increase in PSNR, $13.0\%$ in SSIM, and a $20.8\%$ decrease in LPIPS compared to the leading SUDS [35] baseline. For novel view synthesis, our technique synthesizes high-quality views of unseen timestamps and also significantly surpasses the best EmerNeRF [45] performance by $8.4\%$ in PSNR, $11.3\%$ in SSIM, and $27.3\%$ in LPIPS. These enhancements are visually corroborated in Fig. 4, where our method faithfully captures both static and dynamic elements with remarkable detail. Noteworthy is our method’s efficiency, completing the entire training process around an hour, and stands out as the method with the fastest training speed compared to the baselines. We also report training results of our method on fully 5 cameras, for we use every forth frame of 5 cameras as our testing set, our performance dropped off slightly. Our rendering speed, measured in Frames Per Second (FPS), significantly outperforms competing methods and stands slightly below that of 3DGS [15].
For fair comparison with EmerNeRF [45], we adopt its training setup, i.e., 4$\times$ downsample rate for image resolution and using the whole sequences for training. We randomly selected four Waymo scenes for test in Table 3. It is shown that PVG is clearly superior over EmerNeRF. More visualization is given in Fig. 8.
Table 2: Dynamic scenes’ PSNR, same settings as EmerNeRF [45].
Table 3: Static scenes’ PSNR, same settings as StreetSurf [14].
Sequence Image Reconstruction Novel View Synthesis
EmerNeRF PVG (Ours) EmerNeRF PVG (Ours)
seg102319… 28.93 34.15 28.64 29.52
seg117188… 28.42 32.34 26.29 27.16
seg103913… 29.10 32.01 27.73 28.47
seg109636… 28.39 31.46 27.35 28.28
Average 28.69 32.49 27.50 28.35
Sequence Image Reconstruction
StreetSurf PVG (Ours)
seg153495… 27.26 31.22
seg405841… 28.08 34.91
seg342571… 29.42 32.37
seg134763… 28.20 31.11
Average 28.24 32.40
Table 3: Static scenes’ PSNR, same settings as StreetSurf [14].
(a) S-NeRF [43]
(b) StreetSurf [14]
(c) 3DGS [15]
(d) PVG (Ours)
(e) GT
Figure 6: Novel view synthesis on static scene on Waymo.
For static scenes, as shown in Table 5, we align with the settings in S-NeRF [43], employing all five cameras for training and designating every fourth timestamp’s frame as the test set. The training sequences utilized are the same as those reported in S-NeRF[43]. Our model not only outperforms the baselines [43, 14, 15] which are focused on static scenes across three metrics but also demonstrates a discernible enhancement in image quality, as evidenced in Fig. 6. To be more convinced, we conduct additional experiments using the same setup as in StreetSurf [14], using the randomly selected four sequences from the results reported in StreetSurf[14]. Our evaluation, focusing on PSNR, as StreetSurf only reports PSNR in their papers, reveals in Table 3 that our model achieves significant advancements over previous approaches dedicated to static scenes.
Figure 7: PVG vs. EmerNeRF [45].
PVG Refer to caption
EmerNeRF Refer to caption
GT Refer to caption
RGB Refer to caption
Depth Refer to caption
Semantic Refer to caption
Figure 7: PVG vs. EmerNeRF [45].
Figure 8: Rendered rgb, depth and semantic label.
We also show that by making use of the 2D semantic labels provided in a video sequence, our method can derive semantic categories in novel viewpoints or timestamps and render as shown in Fig. 8. This is achieved by assigning a 19-dimension vector for each Gaussian, representing the probability of which category the Gaussian belongs to and the rendered probability map is supervised by 2D semantic labels with cross-entropy.
Results on KITTI Our approach is also quantitatively evaluated on the KITTI benchmark, following SUDS [35]. We select sequences characterized by extensive movement for analysis. The proposed method surpasses all competitors across every evaluated metric. Notably, while these sequences present a substantial number of dynamic objects, our temporal smoothing mechanism secures a concise scene representation, thereby mitigating over-fitting and ensuring superior image quality in novel viewpoints. Table 1 and Fig. 5 demonstrate our significant improvement over the leading SUDS [35] in image reconstruction, with improvements of $16.0\%$ in PSNR, $7.0\%$ in SSIM, and a $62.2\%$ decrease in LPIPS and surpass the leading EmerNeRF [45] in novel view synthesis with improvement of $8.7\%$ in PSNR, $11.9\%$ in SSIM, and a $51.9\%$ decrease in LPIPS. 3DGS achieves much inferior results on KITTI as there were much more dynamic objects in the scenes than Waymo.
4.2 Ablation study
Table 4: Novel view synthesis of static scene, same settings as SNeRF [43]
Table 5: Ablation on novel view synthesis.
[…]
S-NeRF [43] 23.60 0.743 0.422
StreetSurf [14] 23.08 0.806 0.396
3DGS [15] 24.08 0.807 0.322
PVG (Ours) 26.06 0.832 0.266
[…]
w/o temporal smoothing 26.50 0.805 0.315
w/o depth loss $\mathcal{L}_{d}$ 28.10 0.849 0.279
w/o velocity loss $\mathcal{L}_{\bm{\bar{v}}}$ 28.04 0.847 0.281
w/o sky refinement 28.01 0.846 0.278
w/o position-aware control 27.90 0.846 0.282
PVG (Ours) 28.11 0.849 0.279
Table 5: Ablation on novel view synthesis.
Figure 9: Temporal smoothing (a) on and (b) off.
Refer to caption (a) On Refer to caption (b) Off
Refer to caption (a) Constant Refer to caption (b) Linear Refer to caption (c) Ours
Figure 9: Temporal smoothing (a) on and (b) off.
Figure 10: Dynamics models: (a) Constant, (b) Linear, (c) Ours, with PSNR=27.77/27.09/28.11, respectively.
(a) Effect of the depth $\mathcal{L}_{d}$ and sky module.
(b) Effect of temporal smoothing for motion trend (car, pedestrian and shadow) and sparse velocity $\mathcal{L}_{\bar{v}}$ for enabling static component modeling. We render the average velocity map $\mathcal{\bar{V}}$, and color its uv component by optical flow color coding.
(c) Effect of Positional-Aware Control (PAC) in improving the depth of distant components.
Figure 11: More visualization of the ablation study.
In Table 5, we conduct an ablation study to investigate the impact of the primary components of our model on novel view synthesis within dynamic scenes, specifically utilizing the Waymo Open Dataset. We set $\eta=1$ to deactivate the scene flow-based temporal smoothing mechanism, a crucial element in RGB rendering. As illustrated in Fig. 10, the temporal smoothing mechanism results in enhanced smoothness in novel view rendering and promotes temporal and spatial consistency of PVG points. Meanwhile, temporal smoothing can help the model to learn the correct motion trend as shown in Fig. 10(b). Our findings indicate that the integration of LiDAR supervision, sky refinement module contributes to more plausible geometry (see in Fig. 10(a)). Depth loss and sky module make a minimal effect on novel view RGB rendering mainly because test novel views are interpolations of training views (both in the driving paths), making these metrics not fully reflect the quality of synthesis. Fig. 10(c) demonstrates positional-aware control strategy significantly improves the reconstruction of distant views. Additionally, Fig. 10(b) shows the inclusion of velocity loss facilitates the convergence of velocity to a sparser rank, thereby simplifying the segmentation process between dynamic and static elements.
Dynamics formulation We investigate the influence of $\widetilde{\bm{\mu}}(t)$ formulation. We compare our design (Eq. (6)) with two alternatives: (a) A constant model, i.e., $\widetilde{\bm{\mu}}(t)=\bm{\mu}$, and (b) A linear model, i.e., $\widetilde{\bm{\mu}}(t)=\bm{\mu}+\bm{v}(t-\tau)$. We observe from Fig. 10 that (1) The constant model adeptly represents static elements (see yellow box) but exhibits suboptimal performance on dynamic elements (red box); (2) The linear model can more smoothly capture dynamic aspects. However, it consistently misrepresents static components resulting in ambiguities and ghosting effects. Our method overcomes both limitations by using a sine function based vibration design.
5 Conclusion
We present the Periodic Vibration Gaussian (PVG), a model adept at capturing the diverse characteristics of various objects and materials within dynamic urban scenes in a unified formulation. By integrating periodic vibration, time-dependent opacity decay, and a scene flow-based temporal smoothing mechanism into the 3D Gaussian Splatting technique, we have established that our model significantly outperforms the state-of-the-art methods on the Waymo Open Dataset and KITTI benchmark, with significant efficiency advantage in dynamics scene reconstruction and novel view synthesis. While PVG excels in managing dynamic scenes, it encounters limitations in precise geometric representation, attributable to its highly adaptable design. Future efforts will focus on improving geometric accuracy and further refining the model’s proficiency in accurately depicting the complexities of urban scenes.
Appendix 0.A More model interpretation
Representation
To facilitate understanding how our proposed PVG, we consider a simplified scene with both static and dynamic components, as illustrated in Fig. 12. Concretely, PVG points with long lifespans are used for quantifying static scene elements as the conventional 3D Gaussian counterpart, whilst those with short lifespans distributed over space and time for representing the unconstrained dynamic components of a scene.
Figure 12: Consider a 4D space-time coordinate system with a non-zero slope for dynamic objects and zero slope for the static. Every PVG point’s trajectory is characterized by a piecewise sine function with specific domain of definition and amplitude. PVG points with small staticness coefficient $\rho$ (red points) and short lifespans learn to model dynamic scene parts, alongside those with large $\rho$ (green points) and long lifespans for explaining static scene parts. To represent unconstrained motion (e.g., moving car), a collection of PVG points will work out in a cohort.
Temporal smoothing by intrinsic motion
Due to sparse training frames, the renderings of novel timestamps are prone to underfitting. To make the $\{\mathcal{H}_{i}\}$ temporally and spatially consistent with its intrinsic motion, we exploit the inherent temporal consistency law: The status $\mathcal{H}_{i}(t)$ at time $t$ is similar to the result ^$\widehat{\mathcal{H}}_{i}(t)$ of translating the status to $t-\Delta t$ by a distance $\bm{v}\Delta t$. This introduces an additional optimization regularization defined as:
$\min\mathbb{E}_{t,\Delta t}\mathbf{Diff}(\{\mathcal{H}_{i}(t)\},\{\widehat{\mathcal{H}}_{i}(t)\}),$ (14)
where $\mathbf{Diff}(\cdot)$ is a difference measurement about two set of 3D Gaussian points. While it is hard for us to directly calculate $\mathbf{Diff}(\cdot)$, we use an indirect measurement by rendering function $\mathrm{Render}(\cdot)$.
Our final objective function could be written as the expectation form of the differences between two ways of rendering:
$\min\mathbb{E}_{t,\Delta t,\mathbf{E},\mathbf{I}}\|\mathrm{Render}(\{\mathcal{H}_{i}(t)\})-\mathrm{Render}(\{\widehat{\mathcal{H}}_{i}(t)\})\|,$ (15)
for any camera extrinsic $\mathbf{E}$ and intrinsic $\mathbf{I}$, timestamp $t$ and small time shift $\Delta t$ (the camera pose and time parameter are omitted for simplicity).
To compute Eq. (15), we need to render twice, making sampling every camera pose and timestamp computationally expensive. For efficiency, in practice we sample $\{t,\mathbf{E},\mathbf{I}\}$ uniformly from the training set and sample $\Delta t$ from $U(-\delta,+\delta)$, and replace $\mathrm{Render}(\{\mathcal{H}_{i}(t)\})$ with the ground truth image to only render once for every step. Fig. 13 shows a more vivid illustration of the temporal smoothing mechanism.
Figure 13: Consider two adjacent training frames with timestamp $t_{1}$ and $t_{2}$ (a small time window). In a small time period, we assume dynamic objects move linearly. At observed times $t_{1}$ and $t_{2}$, RGB renderings could fit well. However for the moments in between $(t_{1}<t_{b}<t_{2})$, we have no corresponding training data to constrain our model $\{\mathcal{H}_{i}(t_{b})\}$. To prevent our model from behaving improperly, we impose a smooth constraint subject to the slope of $\bm{v}$. The frames used to train are knots of a function, what we need to do is making the knots more smoothly connected.
Appendix 0.B More implementation details
For experiments on KITTI dataset and whole Waymo sequences compared with EmerNeRF [45], we adjusted some parameters for accommodating more training frames in these data. We train our model for 40000 iterations and densify and prune the gaussians every 200 iterations until 20000 iterations. To better scatter points in 4D space, when we split a PVG point, we shrink each new point’ scales by a decay rate of 0.8 and randomly disturb the $\tau$ with $\Delta\tau$ sampled in $\mathcal{N}(0,\beta^{2})$ and disturb its position with $\Delta\tau\cdot\bar{\bm{v}}$. In the first 10000 iterations, we don’t shrink $\beta$ and in the next 10000 iterations we shrink $\beta$ to $0.8\beta$ in split operation. In clone operation, we just copy every parameter in the PVG point to a new point.
For experiments with StreetSurf [14], we deactivate its pose refinement module for a fairer comparison. Otherwise, the refined poses can not align with the GT poses which leads to a low level of PSNR in the novel view synthesis task.
Appendix 0.C Velocity map and staticness map
(a) render RGB
(b) Velocity map
(c) Staticness map
Figure 14: Visualization of (b) the velocity map and (c) $\rho$ map of (a) a scene with a left-to-right moving car and two walking pedestrians. It is evident that our model captures the motion, dynamic (including even the car’s shadow) and static parts of the scene. In $\rho$ map, blue/red: large/small $\rho$ pointing to static/dynamic areas.
We visualize the renderings of velocity map $\mathcal{\bar{V}}$ and staticness $\rho$ map to analyze the behavior of PVG. For the visualization of $\mathcal{\bar{V}}$, we first transform $\bm{\bar{v}}$ of each pixel from the world coordinate system to the camera coordinate system and project the $\bm{v}_{cam}$ onto a plane parallel to the camera plane. Then we use the color coding of optical flow for visualization.
For the visualization of staticness $\rho$ map, we need to clamp each point’s $\rho$ to the range of $[0,2]$ and render, otherwise the visualization is not visually distinctive.
As shown in Fig. 14, our model captures not only dynamic objects like cars and people, but also moving light and shadow. Note that, pixels with small $\rho$ point to dynamic areas and large $\rho$ for static ones.
Appendix 0.D Visualization for Waymo
More novel view synthesis results are in Fig. 16 and Fig. 17. We project the LiDAR points to the camera world as the ground truth depth map. SUDS [35] is good at image reconstruction, but poor at novel view synthesis. 3DGS [15] is only good for static and close-range reconstruction but unable to handle distance view and dynamic objects. In contrast, our method can not only reconstruct both the near and distance well, but also render images with the quality as the ground-truth.
We also test the rendering quality of PVG under different camera settings, for evaluating the robustness of novel view synthesis. As in Fig. 15, we zoom in/out, disturb the camera extrinsics and intrinsics for novel view synthesis, and PVG remain stable and compact rendering quality. $\mathcal{L}_{d}$ improves geometry and (e.g., road and dynamics) and increases generalization performance beyond the driving path (Fig. 14(d) vs. 14(e)).
(a) Origin
(b) Zoom in
(c) Zoom out
(d) Shift right
(e) w/o $\mathcal{L}_{d}$
Figure 15: Rendered images under different camera settings.
Appendix 0.E Visualization for KITTI
The scenarios of KITTI used in [35] are almost the scenarios where the ego vehicle is not moving. The quality of depth reconstruction depends on the point cloud. Image reconstruction results and novel view synthesis results are in Fig. 18 and Fig. 19.
Appendix 0.F Results on small-scale data
Table 6: Per-scene results on the Plenoptic Video dataset.
Coffee Martini Cook Spinach Cut Beef Flame Salmon Flame Steak Sear Steak Mean
HexPlane [7] - 32.04 32.55 29.47 32.08 32.39 31.70
K-Planes-explicit [12] 28.74 32.19 31.93 28.71 31.80 31.89 30.88
K-Planes-hybrid [12] 29.99 32.60 31.82 30.44 32.38 32.52 31.63
MixVoxels [37] 29.36 31.61 31.30 29.92 31.21 31.43 30.80
NeRFPlayer [29] 31.53 30.56 29.353 31.65 31.93 29.12 30.69
HyperReel [2] 28.37 32.30 32.922 28.26 32.20 32.57 31.10
PVG (ours) 28.64 32.85 33.14 29.65 33.45 33.43 31.86
For further evaluate the capability of our model, we evaluate PVG on a small-scale benchmark Plenoptic Video for dynamic 3D reconstruction. As shown in Table 6, our model outperforms previous methods which are particularly designed for small-scale dynamic reconstruction by a clear margin.
Appendix 0.G Limitation
Unlike neural network based representation models, our PVG uses independent and discrete points to represent a scene. Despite the advantages including high flexibility, simple composition, strong expression ability, and strong fitting ability, the independence of each point makes it more difficult to model the smoothness over time and space. Our temporal smoothing mechanism does enhance the correlation between points, but still not fully solves this problem. More fine-grained, dedicated designs are needed for further enhancement.
S-NeRF [43] Refer to caption
StreetSurf [14] Refer to caption
3DGS [15] Refer to caption
NSG [23] Refer to caption
SUDS [35] Refer to caption
Mars [41] Refer to caption
EmerNeRF [45] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 16: Qualitative results of novel view synthesis on Waymo. GT: Ground-truth.
S-NeRF [43] Refer to caption
StreetSurf [14] Refer to caption
3DGS [15] Refer to caption
NSG [23] Refer to caption
SUDS [35] Refer to caption
Mars [41] Refer to caption
EmerNeRF [45] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 17: Qualitative results of novel view synthesis on Waymo.
S-NeRF [43] Refer to caption
StreetSurf [14] Refer to caption
3DGS [15] Refer to caption
NSG [23] Refer to caption
SUDS [35] Refer to caption
Mars [41] Refer to caption
EmerNeRF [45] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 18: Qualitative results of image reconstruction on KITTI.
S-NeRF [43] Refer to caption
StreetSurf [14] Refer to caption
3DGS [15] Refer to caption
NSG [23] Refer to caption
SUDS [35] Refer to caption
Mars [41] Refer to caption
EmerNeRF [45] Refer to caption
PVG (Ours) Refer to caption
GT Refer to caption
Figure 19: Qualitative results of novel view synthesis on KITTI.